OpenStudy (anonymous):

what is 3/18 written in simplest form? explain how you found your answer

OpenStudy (anonymous):

1/6 3/3=1 18/3=6, put them together for 1/6

OpenStudy (31356):

3/18 3=3*1 18=3*6 (3/3)*(1/6) 3/3=1 1/6

OpenStudy (31356):

Medal?

OpenStudy (31356):

@kval2036